OpenStudy (anonymous):

Theory of equation depressed cubic @ganeshie8

OpenStudy (anonymous):

If a , b , c are the roots of the equation x^3 -x -1 =0 then, \[\frac{ 1+a }{ 1-a }+\frac{ 1+b }{ 1-b }+\frac{ 1+c }{ 1-c }\] Has the value equal to what?

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

Should we consider the x^2 term as 0

OpenStudy (anonymous):

So it would become \[\huge a+b+c = 0\] \[\huge ab+ac+bc = -1\] \[\huge abc=1\]

mathslover (mathslover):

Why would we consider x^2 term as zero? And is there any x^2 term?

OpenStudy (anonymous):

nope it is a depressed one

mathslover (mathslover):

Depressed one? Sorry, can you please explain.

OpenStudy (anonymous):

Depressed means just when there is a term missing in a normal (any) polynomial here x^2 term is missing so it is depressed

mathslover (mathslover):

Oh, okay...

OpenStudy (anonymous):

I am getting thousands of equations but not sure which to use i really need help Or maybe i am making it too complicated somewhat

mathslover (mathslover):

I'm not sure. May be @Hero @whpalmer4 @jim_thompson5910

OpenStudy (anonymous):

do u have an answer for it ? like the finl sol ?

OpenStudy (anonymous):

Maybe this could be of any help: if you just get it to commun denominator, in the first term you get something like this: (1+a)(1-b)(1-c)+.... now, on the other side, since equation has coeficient besides x^2 equal to 0, this means that: (x-a)(x-b)(x-c)=(x-a)(x^2-cx-bx+bc)=0 so putting this two facts together: x^2-cx^2-bx^2=0 or 1-c=b

OpenStudy (anonymous):

The answer is -7

OpenStudy (anonymous):

@thomaster And thank you @myko

OpenStudy (anonymous):

@AkashdeepDeb Its of CBSE help me I am new to this board

OpenStudy (anonymous):

i would try solve for x :P

OpenStudy (anonymous):

so you would have a real , b,c complex lol shall we try that ?

OpenStudy (anonymous):

continuing with my start, :). Sincé the factors (x-a)(x-b)(x-c) can be put in any orther this means that 1-c=a, 1-a=c ... and so on

OpenStudy (anonymous):

Let's seee but i doubt it would lead to something fruitful i already have around 150 sums to complete and most of them are from jee

OpenStudy (anonymous):

Not that i doubt your mathematical ability

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=+x^3+-x+-1+%3D0

OpenStudy (anonymous):

That's cardano's formula , that won't lead anywhere bu to our own doom

OpenStudy (anonymous):

basicly you get 3 equalities: 1-c-b=0 1-c-a=0 1-a-b=0 from here b=a and b+a=1

OpenStudy (anonymous):

so a=b=0.5

OpenStudy (anonymous):

and c you get from any other relation

OpenStudy (anonymous):

Substituting the values i am not getting -7 which is the correct answer @myko

OpenStudy (anonymous):

but the answer would be 9, not -7. So maybe I made a mistake somewhere...:(

OpenStudy (anonymous):

It is comming -1.6666666666666666666................. :-(

OpenStudy (akashdeepdeb):

\[\frac{1+a}{1-a} + \frac{1+b}{1-b} + \frac{1+c}{1-c}\] \[ = \frac{(1+a)(1-b)(1-c) + (1+b)(1-a)(1-c) + (1+c)(1-a)(1-b)}{(1-a)(1-b)(1-c)}\] \[ = \frac{3abc-ab-ac-bc-a-b-c+3}{-abc+ab+ac + bc -a-b-c+1}\] \[=\frac{3abc - (ab+bc+ac) - (a+b+c) + 3}{-abc + (ab+ac+bc) - (a+b+c) + 1}\] Now use the equations that you can get from the cubic equation to get your answer.

OpenStudy (anonymous):

Okay , Are we supposed to include the x^2 term its coefficient being 0 and thank you akshay for being patient and explaing this. :)

OpenStudy (akashdeepdeb):

Yes, the co-efficient of the \(x^2\) HAS to be zero because it is non-existent. So you can write it like this: 0.\(x^2\) And my name is Akashdeep [or Akash] and not Akshay. :)

OpenStudy (anonymous):

Ok AKASHDEEP thank you

OpenStudy (anonymous):

I am wondering what is wrong woth my reasoning? Can`t see the mistake...

OpenStudy (anonymous):

ok found it, :)

OpenStudy (anonymous):

what was it

OpenStudy (anonymous):

a+b+c=0

OpenStudy (anonymous):

but anyway answr is 7 not -7

OpenStudy (anonymous):

i got -7 now it is correct